# Human Mesh Modeling for Anny Body

### v1

@@ 5 Assessing 3D modeling capabilities
To assess how Anny can model the geometry of actual human meshes, we register it to scans of 3DBodyTex [saint20183dbodytex]. 3DBodyTex [saint20183dbodytex] is a dataset consisting of 400 3D scans of adults depicting 100 males and 100 female individuals in minimal clothing. We report qualitative results in Fig. 5. Anny approximates 3D points of the scans with an average point to mesh error of 2.4mm, excluding head and hands of the numerical evaluation due to the low scan quality in these regions, following an existing protocol [osman2022supr, park2025atlas]. Smaller fitting errors have been reported with other human mesh models, notably a 1.8mm mean vertex to vertex error using ATLAS [park2025atlas], however these models are specifically optimized to capture the geometry of individuals wearing minimal clothes. Anny models neither clothes nor the shape deformation they cause, yet is able to model these 3D scans with an accuracy sufficient for many applications. To assess modeling capabilities with children, we register Anny to three commercial scans of children from RenderPeople, and achieve a 2.7mm point to mesh error, excluding regions corresponding to the hairnet worn by the models of the numerical evaluation.
@@ 6 Recovering Anny Human Meshes
Human Mesh Recovery (HMR) is a natural downstream task for assessing the representational power of a body model [multi-hmr2024, patel2025camerahmr]. Recent works [bedlam, tesch2025bedlam2] have shown that large-scale synthetic data can be as effective as, or even superior to, real data—whose annotations often contain inherent noise—for training HMR models. Having a body model interoperable with other meshes that can accommodate clothing enables to easily generate large-scale, diverse synthetic training data. In this paper, we leverage these properties and introduce Anny-One, a synthetic dataset of images with corresponding Anny annotations (Section 6.1). We then use Anny-One to train HMR models, that we evaluate on standard benchmarks (Section 6.2).
@@ 6.1 The Anny-One dataset
For each individual in the dataset, we generate both a ground-truth annotation using Anny and a corresponding HumGen3D character. Each character is randomly augmented with clothing and accessories sampled from the HumGen3D library, ensuring high visual diversity. These characters are then positioned within detailed, procedurally generated indoor scenes created using Infinigen Indoors [infinigen2024indoors]. Scene placement ensures no interpenetration between humans and surrounding objects. On average, each scene includes approximately five people, and we render up to 40 camera views per scene, with camera placement biased toward human-centric framing. We further enhance viewpoint diversity by rendering egocentric perspectives, upper-body close-ups, and hand-focused views for a subset of the data. The camera’s field of view is uniformly sampled between 30° and 130° to capture a wide range of spatial compositions.
Body poses are randomly sampled from AMASS [mahmood2019amass], while hand poses are independently drawn from GRAB[taheri2020grab]. Body shapes are sampled from a statistical distribution of human phenotypes derived from WHO population data (Section 5). To ensure physical plausibility, we apply a self-collision check to eliminate invalid or unrealistic configurations.
All images are rendered in Blender at a resolution of 1280×1280 pixels. Figure 6 shows representative samples from Anny-One, illustrating its diversity in body shapes, clothing, poses, and scene composition.
Overall, Anny-One provides a large-scale, visually rich, and statistically diverse resource for training HMR models in indoor environments.
@@ 6.2 HMR models
Evaluation benchmarks.
We evaluate 3D mesh prediction accuracy on standard benchmarks: 3DPW [vonMarcard2018], EMDB [kaufmann2023emdb], Hi4D [yin2023hi4d], CMU-Toddler [cmu_panoptic_studio] and EHF [pavlakos2019smplx]. We report the commonly used Mean Per Joint Position Error (MPJPE) and Per Vertex Error (PVE), along with their Procrustes-aligned variants, following prior work [conman, multi-hmr2024]. We also report the Pair-PA-MPJPE, which measures the mean joint position error after Procrustes alignment of each pair of interacting humans, thereby evaluating the accuracy of their relative 3D poses. Although the CMU-Toddler dataset contains both adults and children, its limited diversity among children prevents extensive ablation studies. To better assess performance across diverse populations, we additionally use AGORA [agora], which includes both adults and children. While AGORA is synthetic, it remains the only dataset currently available for this purpose. Because the official validation set of AGORA does not contain any children, we re-define the train and val set of AGORA such that images with human scans from 3DPEOPLE constitute the new validation set composed of 2k images, ensuring an equal percentage of children in both training and validation splits.
6.3 HMR Results
@@
Table 2: Effect of Model and Training Data on AGORA [agora] validation set (with diverse shapes from adults and children) using Multi-HMR [multi-hmr2024]. We use either BEDLAM or Anny-One for pre-training and SMPL-X/SMIL-X/Anny as body model.
@@
Anny vs SMPL-X: a simple body model is enough.
In order to compare Anny with SMPL-X [pavlakos2019smplx] in a model-agnostic manner, we adapt the implementations of two foundation models (HMR2.0 [goel2023humans] and Multi-HMR [multi-hmr2024]) to predict full-body meshes using either the Anny or SMPL-X body model. We train the HMR models on BEDLAM and evaluate them on the 3DPW and EHF test sets. To limit the computational cost of experiments, HMR2.0 is finetuned with a frozen backbone and without data augmentation.
Results are shown in Table 1. Across all metrics, Anny achieves comparable or superior performance to SMPL-X, despite its simpler and non data-driven design. This demonstrates that a compact, interpretable, and geometry-consistent body model such as Anny is sufficient for HMR and can serve as a drop-in replacement for SMPL-X.
@@
We then consider training directly on AGORA, and compare Anny to both SMIL-X and SMPL-X. Both SMIL-X and Anny perform significantly better than SMPL-X, with a moderate advantage for Anny. This is consistent with the fact that they were both designed to handle children. Finally, we consider the full-data regime with pretraining on either BEDLAM or Anny-One, followed by finetuning on AGORA (last three rows of the table). We observe that in this large-scale regime, the combination of Anny-One and Anny significantly outperforms existing datasets and body models on AGORA.
@@
We evaluate Multi-HMR trained with Anny on a large-scale mixture of datasets, including Anny and standard HMR training data (BEDLAM, MS-COCO, MPII), and compare against state-of-the-art multi-person HMR approaches. Results on 3DPW, EMDB, Hi4D, and CMU-Toddler (Table 3) show that our model achieves state-of-the-art performance across all benchmarks, consistently outperforming prior methods despite using a simpler, non data-driven, and interpretable body representation, which is of high interest for many downstream applications.
These results demonstrate that Anny scales effectively to large and diverse data, enabling robust reconstruction of adults and children alike across both in-the-wild and controlled multi-person scenarios. Anny achieves this performance without relying on data-driven models demonstrating that a simple, unified and interpretable human body model can scale effectively to modern HMR tasks.
@@ 7 Conclusion
We present Anny, a unified, differentiable, interpretable, and non data-driven human body model that continuously encode human shape variation across age, gender, and body type. Built entirely on open and procedural knowledge, Anny bridges the gap between artistic and statistical modeling, offering a transparent alternative to abstract latent spaces. Despite not being learned from 3D scans, Anny achieves competitive performance in fitting real scans, showing that accurate geometric modeling can be obtained purely from procedural priors and anthropometric calibration. Using Anny, we generate Anny-One, a large-scale dataset of 800k synthetic images with rich diversity in pose, shape, and scene context. Experiments show that HMR models trained with Anny achieve state-of-the-art performance across diverse benchmarks. By providing a free, interpretable and unified 3D model that covers the full human lifespan, we hope Anny will broaden the ecosystem of human-aware computer vision.

### v2

@@ 5 Assessing 3D modeling capabilities
To assess how Anny can model the geometry of actual human meshes, we register it to scans of 3DBodyTex [saint20183dbodytex]. 3DBodyTex is a dataset consisting of 400 3D scans of adults depicting 100 male and 100 female individuals in minimal clothing. We report qualitative results in Fig. 5. Anny approximates 3D points of the scans with an average point to mesh error of 2.4mm, excluding head and hands of the numerical evaluation due to the low scan quality in these regions, following an existing protocol [osman2022supr, park2025atlas]. Smaller fitting errors have been reported with other human mesh models, notably a 1.8mm mean vertex to vertex error using ATLAS [park2025atlas], however these models are specifically optimized to capture the geometry of individuals wearing minimal clothes. Anny models neither clothes nor the shape deformation they cause, yet is able to approximate 3D scans with an accuracy sufficient for many applications. To assess modeling capabilities with children, we register Anny to three commercial scans of children from RenderPeople, and achieve a 2.7mm point to mesh error, excluding regions corresponding to the hairnet worn by the models of the numerical evaluation.
@@ 6 Recovering Anny Human Meshes
Human Mesh Recovery (HMR) is a natural downstream task for assessing the representational power of a body model [multi-hmr2024, patel2025camerahmr]. Recent works [bedlam, tesch2025bedlam2] have shown that large-scale synthetic data can be as effective as, or even superior to, real data—whose annotations often contain inherent noise—for training HMR models. Having a body model interoperable with other models that can accommodate clothing enables to easily generate large-scale, diverse synthetic training data. In this paper, we leverage these properties and introduce Anny-One, a synthetic dataset of images with corresponding Anny annotations (Section 6.1). We then use Anny-One to train HMR models, that we evaluate on standard benchmarks (Section 6.2).
@@ 6.1 The Anny-One dataset
For each individual in the dataset, we generate both a ground-truth annotation using Anny and a corresponding HumGen3D character. Each character is randomly augmented with clothing and accessories sampled from the HumGen3D library, ensuring high visual diversity. These characters are then positioned within detailed, procedurally generated indoor scenes created using Infinigen Indoors [infinigen2024indoors]. Scene placement ensures no interpenetration between humans and surrounding objects. On average, each scene includes approximately five people, and we render up to 40 camera views per scene, with camera placement biased toward human-centric framing. We further enhance viewpoint diversity by rendering egocentric viewpoints, upper-body close-ups, and hand-focused views for a subset of the data. The camera’s field of view is uniformly sampled between 30° and 130° to capture a wide range of spatial compositions. Body poses are randomly sampled from AMASS [mahmood2019amass], while hand poses are independently drawn from GRAB [taheri2020grab]. Body shapes are sampled from a statistical distribution of human phenotypes derived from WHO population data (Section 5). To ensure physical plausibility, we apply a self-collision check to eliminate invalid or unrealistic configurations. All images are rendered using Blender Cycles [blender] at a resolution of 1280×1280 pixels. Figure 6 shows representative samples from Anny-One, illustrating its diversity in body shapes, clothing, poses, and scene composition. Overall, Anny-One provides a large-scale, visually rich, and statistically diverse resource for training HMR models in indoor environments.
@@ 6.2 HMR models
Table 2: Effect of Model and Training Data on AGORA [agora] validation set (with diverse shapes from adults and children) using Multi-HMR [multi-hmr2024]. We use either BEDLAM or Anny-One for pre-training and SMPL-X/SMPL-X+A/Anny as body model.
Evaluation benchmarks.
We evaluate 3D mesh prediction accuracy on standard benchmarks: 3DPW [vonMarcard2018], EMDB [kaufmann2023emdb], Hi4D [yin2023hi4d], CMU-Toddler [cmu_panoptic_studio] and EHF [pavlakos2019smplx]. We report the commonly used Mean Per Joint Position Error (MPJPE) and Per Vertex Error (PVE), along with their Procrustes-aligned variants, following prior work [conman, multi-hmr2024]. We also report the Pair-PA-MPJPE, which measures the mean joint position error after Procrustes alignment of each pair of interacting humans, thereby evaluating the accuracy of their relative 3D poses. Although the CMU-Toddler dataset contains both adults and children, its limited diversity among children prevents extensive ablation studies. To better assess performance across diverse populations, we use AGORA [agora], which includes both adults and children. While AGORA is synthetic, it remains the only dataset currently available for this purpose. Because the official validation set of AGORA does not contain any children, we re-define the training and validation sets such that images with human scans from 3DPEOPLE constitute the new validation set composed of 2k images, ensuring an equal percentage of children in both training and validation splits.
Figure 7: Runtime performance of a body model forward pass, measured on a NVIDIA A100 GPU.
6.3 HMR Results
@@
Anny vs. SMPL-X: a scan-free body model is enough.
We adapt the implementations of two foundation models (HMR2.0 [goel2023humans] and Multi-HMR [multi-hmr2024]) to predict full-body meshes using either Anny or the SMPL-X [pavlakos2019smplx] body model. We train the HMR methods on BEDLAM and evaluate them on the 3DPW and EHF test sets. To limit the computational cost of experiments, HMR2.0 is finetuned with a frozen backbone and without data augmentation. We also restrict the modeling of Anny shapes to 6 main phenotype parameters (age, gender, height, weight, muscle, proportions).
Results are reported in Table 1. Across all metrics, using Anny achieves comparable or superior performance to SMPL-X, despite its non data-driven design. This demonstrates that a compact, interpretable, and geometry-consistent body model such as Anny is sufficient for HMR and can serve as a drop-in replacement for SMPL-X.
@@
We then consider training directly on AGORA, and compare Anny to both SMPL-X+A and SMPL-X. Both SMPL-X+A and Anny perform significantly better than SMPL-X, with a moderate advantage for Anny. This is consistent with the fact that they were both designed to handle children. Finally, we consider the full-data regime with pretraining on either BEDLAM or Anny-One, followed by finetuning on AGORA (last three rows of the table). We observe that in this large-scale regime, the combination of Anny-One and Anny significantly outperforms existing datasets and body models on AGORA.
@@
We evaluate Multi-HMR trained with Anny on a large-scale mixture of datasets, including Anny and standard HMR training data (BEDLAM, MS-COCO, MPII), and compare against state-of-the-art multi-person HMR approaches. Results on 3DPW, EMDB, Hi4D, and CMU-Toddler (Table 3) show that our model achieves state-of-the-art performance across all benchmarks, consistently outperforming prior methods.
These results demonstrate that Anny scales effectively to large and diverse data, enabling robust reconstruction of adults and children alike across both in-the-wild and controlled multi-person scenarios. HMR methods can achieve state-of-the-art performance without relying on a scan-based model, demonstrating that a simple, unified and interpretable human body model like Anny can scale effectively to modern HMR tasks.
@@
7 Runtime Performance
Lastly, we plot the runtime performance of Anny in Fig. 7. We experiment using a NVIDIA A100 GPU, and report the performance of previous models [loper2015smpl, pavlakos2019smplx] for reference, using the implementation of V. Choutas. When considering all 267 phenotype parameters, our implementation of Anny shows some runtime overhead compared to SMPL and SMPL-X with small batch sizes (e.g. 49ms for Anny vs. 4ms for SMPL-X when modeling a single person). Anny scales however better and requires significantly less memory for larger batch sizes (e.g. 3GB for Anny vs. 11GB for SMPL-X with a batch size of 8192) thanks to our Warp implementation. Significant speed-up and memory gains can additionally be achieved by restricting the number of phenotype parameters considered – e.g. a 14ms runtime when using 6 phenotype parameters, as we do in our HMR experiments – and by considering a coarser model.
@@ 8 Conclusion
We present Anny, a unified, differentiable, interpretable, and scan-free human body model that continuously encodes human shape variations across age, gender, and body type. Built entirely on open and procedural knowledge, Anny bridges the gap between artistic and statistical modeling, offering a transparent alternative to abstract latent spaces. Despite not being learned from 3D scans, Anny achieves competitive performance in fitting real scans, showing that accurate geometric modeling can be obtained purely from procedural priors and anthropometric calibration. Using Anny, we generate Anny-One, a large-scale dataset of 800k synthetic images with rich diversity in pose, shape, and scene context. Experiments show that HMR models trained with Anny achieve state-of-the-art performance across diverse benchmarks. By providing a free, interpretable and unified 3D model that covers the full human lifespan, we hope Anny will broaden the ecosystem of human-aware computer vision.